Find the length of XW. Round to the nearest hundredth. 180-108 .ZT 854 2) 108 x 360 12 AT= 24 meters meters XW
Answer #1

\angleXZW = 180 - 108 = 72o. (Sum of angles on a straight line = 180o.)

Circumference = 2\pir.

r = 24 / 2 = 12

Therefore Circumference = 2 * \pi * 12 = 24\pi

Therefore Arc XW = (72 / 360) * 2\pir. = (72 / 360) * 24\pi. = 4.8\pi = 4.8 * 22 / 7 = 15.09 meters